OpenStudy (anonymous):

I'm a little stumped with trying to find the cosine equation with these properties: axis of curve y=5 amplitude of 1/4 period of 4π phase shift of π/5 so it would be something like 1/4cos [ ( ) ]+5 ?

OpenStudy (anonymous):

I think I accidentally closed this question on a previous post, so this is a repost of the question, now that I know how to navigate this site, haha.

OpenStudy (anonymous):

\[y = y_1 + Acos(\frac{2 \pi x}{T} + \epsilon)\]Does this help? I won't spell out what all of the constants are for you since you're almost there already, but if you don't see what to do then just ask =)

OpenStudy (anonymous):

Oh, oh no! :( I don't understand. the equation is asking for a y= Acos [K(x-D)] +/- C format? I just don't know how to write in the period and phase shift part

OpenStudy (anonymous):

In the version I gave, I was using T for the period and epsilon (small curved E) for the phase shift. Does that help?

OpenStudy (anonymous):

Uhm, kind of? I'm sorry, haha. So it would be\[\frac{ 1 }{ 4 } \cos [ 5 ( \frac{ 2π }{ 4 } ) ] +5\] ?

OpenStudy (anonymous):

Not quite. I'm not sure how you get that, sorry. Let's just look at the bit inside the brackets; are you going from the equation I gave?